Genuine Spice Inc. began operations on January 1 of the current year. The company produces eight-ounce bottles of hand and boDIRECT LABOR Department Cost Behavior Time per Case Labor Rate per Hour Cost per Case Mixing Variable 20 min. $18.00 $6.00 FiPart B - August Budgets During July of the current year, the management of Genuine Spice Inc. asked the controller to prepareThere was negligible work in process inventory assumed for either the beginning or end of the month; thus, none was assumed.Filling 14.00 5.60 min. Actual variable overhead $305.00 Normal volume 1,600 cases The prices of the materials were differentAmount Descriptions-Part A Controllable variance Equipment depreciation Facility lease Supplies Utilities Volume variance1. Determine the fixed and variable portion of the utility cost using the high-low method. At High Point At Low Point VariablTotal cost 2. Determine the contribution margin per case. $ 3. Determine the fixed costs per month, including the utility fix5. Prepare the August production budget. Enter all amounts as positive numbers. Genuine Spice Inc. Production Budget For theFactory Overhead Volume Variance Normal volume (cases) Actual volume (cases) Difference Factory overhead volume variance The
Answer #1

Part A

1.

At the High Point At the Low Point
Variable cost per unit $                       0.20 $                      0.20
Total fixed cost $                        500 $                       500
Total cost $                        740 $                       600

Working:

Variable utility cost = ($740 - $600)/(1200 - 500) = $140/700 = $0.20 per case

Fixed cost (at the high point) = $740 - ($0.20 x 1200) = $740 - $240 = $500

Fixed cost (at the low point) = $600 - ($0.20 x 500) = $600 - $100 = $500

2. Contribution margin per case: $55.60

Sales price per case $ 100.00
Less: Variable costs
Direct materials 17.00
Direct labor 7.20
Factory overhead (Utilities) 0.20
Selling commission 20.00
Total variable costs per case $ 44.40
Contribution margin per case $ 55.60

3.

1 Total fixed costs:
2 Utilities 500
3 Factory lease 14000
4 Equipment depreciation 4300
5 Supplies 660
6 Total fixed costs $ 19460

4. Break-even number of cases per month: 350

Break-even (units) = Total fixed costs/Contribution margin per unit = $19460/$55.60 = 350 cases

Per HOMEWORKLIB RULES, the first 4 sub-parts have been answered. Please post the remaining separately. Thank you.